OpenStudy (anonymous):

use the elimination method to solve the system -x+2y=10 -3x+6y=11 Options A-infinite solutions B-(-5,2) C-(5,-2) D- no solutions Please show work

OpenStudy (anonymous):

just solve equation one for x x=2y-10 put this in the second equation at the place of x -3(2y-10)+6y=11 so -6y+10+6y=11 can you find that 0y=11 what do you think it is possible?

OpenStudy (anonymous):

wat do u mean

OpenStudy (anonymous):

i mean y=11/0 ?? what does it mean ?

OpenStudy (anonymous):

that does not work that's not even one of the answers

OpenStudy (anonymous):

you have calculator ?

OpenStudy (anonymous):

yes i have the calculator in my hand

OpenStudy (anonymous):

ok just find this and let me know what it says ! 11/0

OpenStudy (anonymous):

ther is no way i could put that in a calculator if 11/0 is 0 duh

OpenStudy (anonymous):

always remember that division by zero is undefined and is not possible!

OpenStudy (anonymous):

yea but i dont get where u got the 11 from its so confusing

OpenStudy (anonymous):

and u cant put the first problem into the second one that is not possible

OpenStudy (anonymous):

look at your second equation there is 11 on the right . and sorry there is a typo in my response! just solve equation one for x x=2y-10 put this in the second equation at the place of x -3(2y-10)+6y=11 so -6y+10+6y=11 so 0y=1 this equation mans means \[\Large 0=1\] which is not possible so system has no solution.

OpenStudy (anonymous):

hes right the answer is D

OpenStudy (anonymous):

hhhhhhhhhhheeeeeeeeellllllllllpppppppppp me still lost

OpenStudy (anonymous):

hhhhhhhhhhhhhhhhheeeeeeeeeeeeeeeeeeeeeeeeeelllllllllllllllllllllllllllpppppppppppppppp

OpenStudy (anonymous):

ok tell me what you are unable to understand. just tell me is it possible that? \[\Large 0=1\]